# The polynomial set associated with a fixed number of matrix-matrix multiplications

### v1

@@ 3.1 Substitution transformations
The following theorems describe transformations of the evaluation scheme $(A,B,c)$ in a way that the output polynomial $p(X)$ is unchanged. The modified evaluation scheme will be denote $(\hat{A},\hat{B},\hat{c})$ and the corresponding $Q$-matrices will be denoted $\hat{Q}_{1},\ldots,\hat{Q}_{m+1}$. The first transformation follows from a change of variables, where we rescale one of the $Q$-coefficients. For example if we set
@@
This implies that without changing the polynomial $p$, we may rescale one row in $A$ and one column in $A$ and $B$. This is formalized in the following theorem, where we omit index on elements of $A$ and $B$ that are unchanged in $\hat{A}$ and $\hat{B}$. From symmetry, we note that the same result holds for $A$ and $B$ switched.
Theorem 3.1.
Let $p$ be the polynomial associated with $(A,B,c)\in\mathbb{C}^{m\times(m+1)}\times\mathbb{C}^{m\times(m+1)}\times\mathbb{C}^{m+2}$ and $\hat{p}$ be the polynomial associated with $(\hat{A},\hat{B},\hat{c})$, where
@@
(14a) $\hat{a}_{k,j}$ $=$ $\alpha a_{k,j},\;\;\;\qquad j=1,\ldots,k+1$
(14b) $\hat{a}_{i,k+2}$ $=$ $\alpha^{-1}a_{i,k+2},\;\;\;i=k+1,\ldots,m$
@@
for any $k\in\left[1,\ldots,m\right]$ and $\0\alpha\in\mathbb{C}\backslash\{0\}$.
Proof 3.2.
@@
To prove (16a), we observe that since the first $k-1$ rows of $\hat{A}$ and $\hat{B}$ are unchanged, the variables $\hat{Q}_{1},\ldots,\hat{Q}_{k+1}$ are also unchanged.
@@
$\hat{Q}_{k+2}$ $=(\alpha a_{k,1}Q_{1}+\cdots+\alpha$ $a_{k,k+1}Q_{k+1}$ $)(b_{k,1}Q_{1}+\cdots+b_{k,k+1}Q_{k+1}$ $)=\alpha Q_{k+2}$
@@
$\hat{Q}_{k+3}$ $=(a_{k+1,1}Q_{1}+\ldots+\alpha^{-1}a_{k+1,k+2}\hat{Q}_{k+2}$ $)(b_{k+1,1}Q_{1}+\ldots+\alpha^{-1}b_{k+1,k+2}\hat{Q}_{k+2}$ )))
@@
where the last equality follows from (16b). The general statement follows from induction.
In a similar fashion we carry out a change of variables where we add one given value $\alpha$ to one of the elements in the first column. For example, adding $\alpha$ to the coefficient corresponding to $Q_{1}=I$, in the first multiplication, leads to
$\hat{Q}_{3}=\left((a_{11}+\alpha)Q_{1}+a_{12}Q_{2}\right)\left(b_{11}Q_{1}+b_{12}Q_{2}\right)$
@@
Hence, a modification in the coefficient corresponding to $Q_{1}$ can be compensated by modifying coefficients in rows below the modification. This can be applied to an arbitrary row $k$ and arbitrary $\alpha$.
Theorem 3.3.
Let $p$ be the polynomial associated with $(A,B,c)\in\mathbb{C}^{m\times(m+1)}\times\mathbb{C}^{m\times(m+1)}\times\mathbb{C}^{m+2}$ and $\hat{p}$ be the polynomial associated with $(\hat{A},\hat{B},\hat{c})$ where
@@
Proof 3.4.
Similarly to the proof of Theorem 3.1, we establish the following three statements:
@@
Together with the definition of $\hat{c}$ in (21c), this implies that the conclusion of the theorem in equation (22) holds. Relation (23a) follows analogously to the proof in Theorem 3.1.
@@
$\hat{Q}_{k+2}$ $=$ $((a_{k,1}+\alpha)Q_{1}+a_{k,2}Q_{2}+\cdots+a_{k,k+1}Q_{k+1})B_{k+2}$
@@
To prove (23c), we prove $\hat{A}_{k+3}=A_{k+3},\ldots,\hat{A}_{m+2}=A_{m+2}$, $\hat{B}_{k+3}=B_{k+3},\ldots,\hat{B}_{m+2}=B_{m+2}$, i.e., that the factors for $\hat{Q}_{k+3}$,…, $\hat{Q}_{m+2}$ are unchanged. For $k+3$ we have
(24) $\hat{A}_{k+3}$ $=$ $(\hat{a}_{k+1,1}Q_{1}+\cdots+\hat{a}_{k+1,k+1}Q_{k+1}+a_{k+1,k+2}\hat{Q}_{k+2})$
@@
where we have inserted (21a) and (23b) to obtain the second equality. The relation $\hat{B}_{k+3}=B_{k+3}$ can be shown analogously using (21b) and (23b). With induction we can prove the corresponding relation for any factor. Consequently, we have
(25) $\hat{Q}_{i+2}=\hat{A}_{i+2}\hat{B}_{i+2}=A_{i+2}B_{i+2}=Q_{k+2},\quad i=k+1,\ldots,m$
@@
The conclusion (22) follows from the same construction as (24).
@@ 3.2 Normalized forms and unreduced schemes
The theorems in the previous section can be applied to impose a certain structure on the triplet $(A,B,c)$ without (or with very little) loss of generality. For any triplet $(A,B,c)$ we can invoke Theorem 3.3 repeatedly. By selecting $\alpha$ as the negative element in the first column, we obtain matrices $A$ and $B$ with a first column containing only zeros. The first column corresponds to the addition of a scaled identity, which is independent (constant) with respect to the input $X$. Note that this can be done for any triplet, and no generality is lost by assuming the first column is zero.
Definition 3.5.
@@
Similarly, if we assume that the Hessenberg matrices $A$ and $B$ are unreduced [11, pg 381], i.e., the elements $a_{1,2},\ldots,a_{m,m+1}$, $b_{1,2},\ldots,b_{m,m+1}$ are non-zero, we can impose further structure by repeatedly applying Theorem 3.1 with scaling determined by the corresponding last non-zero element of the row. This process imposes a normalization on each row.
Definition 3.7.
A constant-free evaluation scheme satisfying $a_{1,2}=\cdots=a_{m,m+1}=b_{1,2}=\cdots=b_{m,m+1}=1$, is called a normalized evaluation scheme and we call the triplet $(A,B,c)$ normalized.
Corollary 3.8.
Any evaluation scheme corresponding to $(A,B,c)$ where $A$ and $B$ are unreduced Hessenberg matrices, is equivalent to a normalized evaluation scheme.
@@ 3.3 Further transformations
For normalized constant-free evaluation schemes, we have $a_{11}=b_{11}=0$ and $a_{12}=b_{12}=1$, meaning that the first multiplication always corresponds to squaring of the input matrix, i.e.,
@@
This assumption can be made without loss of generality and will be used henceforth. This fact was already observed by Paterson and Stockmeyer [19, pg 61].
We now derive further transformations under a technical assumption. For the first two rows we assume that we have the structure of a unreduced constant-free triplet. [^1] This is assumption is made mostly simplify the derivation, and similar results hold, e.g. when $a_{23}=0$. This in turn is equivalent to assuming $a_{23}=b_{23}=1$.
Consider the following transformation based on modifying the second row. We add $\alpha$ and $\beta$ to the second element in the second row in $A$ and $B$ respectively, i.e.,
@@
If we enforce $\beta=-\alpha$, one term in the modification disappears and we can compensate for the other term in a way such that the $\hat{Q}_{5}=Q_{5},\ldots,\hat{Q}_{m}=Q_{m}$ are unmodified.
Theorem 3.9.
Let $p$ be the polynomial associated with a constant-free triplet $(A,B,c)\in\mathbb{C}^{m\times(m+1)}\times\mathbb{C}^{m\times(m+1)}\times\mathbb{C}^{m+2}$, satisfying $a_{23}=b_{23}=1$, and $\hat{p}$ be the polynomial associated with $(\hat{A},\hat{B},\hat{c})$ where
@@
To prove (32a), we use the relation (27). More precisely, we substitute $\beta=-\alpha$ into (27b) and obtain
@@
To prove (32b), we prove that the constructing factors are unchanged, i.e., $\hat{A}_{3}=A_{3},\ldots,\hat{A}_{m+2}=A_{m+2}$ and $\hat{B}_{3}=B_{3},\ldots,\hat{B}_{m+2}=B_{m+2}$. For the first factor equality we have
$\hat{A}_{5}$ $=$ $a_{32}Q_{2}+\hat{a}_{33}Q_{3}+a_{34}\hat{Q}_{4}$
$=$ $a_{32}Q_{2}+a_{33}Q_{3}-a_{34}sQ_{3}+a_{34}Q_{4}+a_{34}sQ_{3}=A_{5},$
where we have used (30a) and (32a) in the second equality. The relation $\hat{B}_{5}=B_{5}$ can be shown analogously using (30b) and (32a). With induction we can prove the corresponding relation for any factor. Consequently, we have
(34) $\hat{Q}_{i+2}=\hat{A}_{i+2}\hat{B}_{i+2}=A_{i+2}B_{i+2}=Q_{i+2},\quad i=3,\ldots,m,$
which proves (32b). The theorem conclusion (31) follows by the same construction as (3.10).
For a given evaluation scheme satisfying $a_{23}=b_{23}=1$, we can apply the previous theorem with $\alpha=b_{22}$, so that $\hat{b}_{22}=0$. This shows that we can assume $b_{22}=0$ for evaluation schemes that are unreduced in the first two rows, without loss of generality. Under these assumptions we get for the second multiplication:
@@
Next, we state and prove a theorem for the third row of the coefficient matrices. For this it is necessary to assume that the first three rows are unreduced, i.e., $a_{23}=a_{34}=b_{23}=b_{34}=1$.
Before proceeding, we illustrate the main idea. The theorem is based on perturbing each of the free elements in the third row of $A$ and $B$, while simultaneously preserving the final output polynomial. In particular, we use perturbations with the following structure:
@@
It turns out that when the perturbations have this particular structure, $Q_{5}$ is modified by the addition of a linear combination of $Q_{3}$, $Q_{4}$ and $X^{3}$. This is helpful because we can compensate for $Q_{3}$ and $Q_{4}$, since we have access to these matrices directly. We can however not compensate for $X^{3}$ directly. Fortunately, we can ensure that the $X^{3}$-coefficient modification is zero by placing an additional condition on $\alpha$ and $\beta$, encoded in the function $z(\alpha,\beta)=0$.
Theorem 3.11.
Let $p$ be the polynomial associated with the constant-free triplet $(A,B,c)\in\mathbb{C}^{m\times(m+1)}\times\mathbb{C}^{m\times(m+1)}\times\mathbb{C}^{m+2}$ satisfying $a_{23}=a_{34}=b_{23}=b_{34}=1$, and $\hat{p}$ be the polynomial associated with $(\hat{A},\hat{B},\hat{c})$ where
@@
where,
(40a) $s_{1}$ $=$ $\alpha(b_{32}-a_{32})-\alpha^{2}$
@@
To prove (43a) we will manipulate the expression for $\hat{Q}_{5}$ until we have $Q_{5}$ plus some remainder term. The goal is to show that this remainder term is a linear combination of $Q_{3}$, $Q_{4}$, and $X^{3}$, where the $X^{3}$-coefficient is given by $z(\alpha,\beta)$, which is zero by assumption. If we express $\hat{Q}_{5}$ in terms of the constructing multiplication factors, we get
@@
From this we can expand and use that $Q_{5}=A_{5}B_{5}$ to get
(45) $\begin{split}\hat{Q}_{5}&=Q_{5}+(B_{5}-A_{5})(\alpha Q_{2}+\beta Q_{3})-(\alpha Q_{2}+\beta Q_{3})^{2},\end{split}$
which is on the desired form. The next step is to show that the remainder is a linear combination of $Q_{3}$, $Q_{4}$, and $X^{3}$. We factorize the remainder in (45) to obtain
@@
(47) $\begin{split}d_{32}=b_{32}-a_{32},\\ d_{33}=b_{33}-a_{33}.\end{split}$
@@
We substitute this expression into (46) and simplify the first factor of the remainder
(49) $\begin{split}\hat{Q}_{5}&=Q_{5}+\big{(}d_{32}Q_{2}+d_{33}Q_{3}-\alpha Q_{2}-\beta Q_{3}\big{)}\big{(}\alpha Q_{2}+\beta Q_{3}\big{)}\\ &=Q_{5}+\big{(}(d_{32}-\alpha)Q_{2}+(d_{33}-\beta)Q_{3}\big{)}\big{(}\alpha Q_{2}+\beta Q_{3}\big{)}.\\ \end{split}$
Next, we expand the remainder
(50) $\begin{split}\hat{Q}_{5}&=Q_{5}+\alpha(d_{32}-\alpha)Q_{2}^{2}+\big{(}\alpha(d_{33}-\beta)+\beta(d_{32}-\alpha)\big{)}Q_{2}Q_{3}+\beta(d_{33}-\beta)Q_{3}^{2}\\ &=Q_{5}+s_{1}Q_{2}^{2}+(\alpha d_{33}+\beta d_{32}-2\alpha\beta)Q_{2}Q_{3}+s_{2}Q_{3}^{2},\\ \end{split}$
@@
where we have used that $z(\alpha,\beta)=0$ in the last equality. This proves the statement.
To prove (43b) we show that the constructing factors are unchanged, i.e., $\hat{A}_{6}=A_{6},\ldots\hat{A}_{m+2}=A_{m+2}$ and $\hat{B}_{6}=B_{6},\ldots\hat{B}_{m+2}=B_{m+2}$. For the first factor equality we have
(53) $\begin{split}\hat{A}_{6}&=(a_{42}Q_{2}+\hat{a}_{43}Q_{3}+\hat{a}_{44}Q_{4}+a_{45}\hat{Q}_{5})\\ &=a_{42}Q_{2}+(a_{43}-a_{45}s_{1})Q_{3}+(a_{44}-a_{45}s_{2})Q_{4}+a_{45}Q_{5}+a_{45}(s_{1}Q_{3}+s_{2}Q_{4})\\ &=A_{6}-a_{45}s_{1}Q_{2}-a_{45}s_{2}Q_{4}+a_{45}(s_{1}Q_{3}+s_{2}Q_{4})\\ &=A_{6}.\end{split}$
where we have used (39a), (39b) and (43a) in the second equality. Note that although $a_{45}=1$, writing the coefficient explicitly better demonstrates how to generalize for $\hat{A}_{i+2}=A_{i+2}$, $i>4$, since $a_{i5}\neq 1$ in general. The relation $\hat{B}_{6}=B_{6}$ follows analogously using (39c), (39d) and (43a).
@@
(54) $Q_{i+2}=\hat{A}_{i+2}\hat{B}_{i+2}=A_{i+2}B_{i+2}=Q_{i+2},\quad i=4,\ldots,m.$
which proves (43b). The theorem conclusion (41) follows by the same construction as (53).
Free variables in Theorem 3.11
Note that Theorem 3.11 includes a scalar valued condition $z(\alpha,\beta)=0$, involving two scalar variables. Let $d_{32}=b_{32}-a_{32}$ and $d_{33}=b_{33}-a_{33}$ be defined as in the proof of the theorem. If we let $\alpha$ be given, we get a quadratic equation in $\beta$:
@@
When $a_{22}\neq 0$, the solution $\beta$ to the equation is explicitly availbable from the solution of the quadratic equation. This has a disadvantage of introducing a square root operation. In a real setting, this can yield complex coefficients.
The paper [24] suggest several approaches to evaluate polynomials with a low number multiplications. For the case $m=3$, the formulas [24, equation (31)] involve a square root, and indeed that approach can be derived from the above transformation with $\alpha=-a_{32}$ and solving for $\beta$.
Suppose $\beta$ is given. Then, the solution for $\alpha$ can be expressed as:
@@
with the condition $\beta\neq\frac{1}{2}d_{33}$. To reframe this condition in terms of entries in matrices $A$ and $B$, we use a change of variables that essentially generalizes [21] where it is given for $m=3$:
@@
This leads to the relation:
@@
Consequently, any evaluation scheme that is unreduced in the first three rows can be transformed into an equivalent scheme satisfying:
@@
as we can choose any non-zero $r$. Thus, when considering evaluation schemes that are unreduced in the first three rows, we can assume without loss of generality that $b_{33}=a_{33}+1$. This assumption helps reduce the number of variables without introducing additional constraints.

### v3

@@ 3.1 Substitution transformations
[…]
@@
This implies that without changing the polynomial $p$, we may rescale one row in $A$ and one column in $A$ and $B$. This is formalized in the following theorem, where we omit indices on elements of $A$ and $B$ that are unchanged in $\hat{A}$ and $\hat{B}$. By symmetry, we note that the same result holds for $A$ and $B$ switched.
Theorem 1.
[…]
@@
(9a) $\hat{a}_{k,j}$ $=$ $\alpha a_{k,j},\;\;\;\qquad j=1,\ldots,k+1,$
(9b) $\hat{a}_{i,k+2}$ $=$ $\alpha^{-1}a_{i,k+2},\;\;\;i=k+1,\ldots,m,$
@@
[…]
Proof.
@@
To prove (11a), we observe that since the first k1k-1 rows of $\hat{A}$ and $\hat{B}$ are unchanged, the variables $\hat{Q}_{1},\ldots,\hat{Q}_{k+1}$ are also unchanged.
@@
$\hat{Q}_{k+2}$ $=(\alpha a_{k,1}Q_{1}+\cdots+\alpha$ $a_{k,k+1}Q_{k+1}$ $)(b_{k,1}Q_{1}+\cdots+b_{k,k+1}Q_{k+1}$ $)=\alpha Q_{k+2},$
@@
$\hat{Q}_{k+3}$ $=(a_{k+1,1}Q_{1}+\ldots+\alpha^{-1}a_{k+1,k+2}\hat{Q}_{k+2}$ $)(b_{k+1,1}Q_{1}+\ldots+\alpha^{-1}b_{k+1,k+2}\hat{Q}_{k+2}$ $)$
@@
where the last equality follows from (11b). The general statement (11c) follows from induction. ∎
In a similar fashion, we carry out a change of variables where we add a given value $\alpha$ to one of the elements in the first column. For example, adding $\alpha$ to the coefficient corresponding to $Q_{1}=I$ in the first multiplication leads to
$\hat{Q}_{3}=\left((a_{11}+\alpha)Q_{1}+a_{12}Q_{2}\right)\left(b_{11}Q_{1}+b_{12}Q_{2}\right),$
@@
Hence, a modification in the coefficient corresponding to $Q_{1}$ can be compensated for by modifying the coefficients in all rows below the modification. This can be applied to an arbitrary row $k$ and arbitrary $\alpha$.
Theorem 2.
[…]
@@
Proof.
Similarly to the proof of Theorem 1, we establish the following three statements:
@@
Together with the definition of $\hat{c}$ in (16c), this implies that the conclusion of the theorem in equation (17) holds. Relation (18a) follows analogously to the proof in Theorem 1.
@@
$\hat{Q}_{k+2}$ $=$ $\left((a_{k,1}+\alpha)Q_{1}+a_{k,2}Q_{2}+\cdots+a_{k,k+1}Q_{k+1}\right)B_{k+2}$
@@
To prove (18c), we prove that $\hat{A}_{k+3}=A_{k+3},\ldots,\hat{A}_{m+2}=A_{m+2}$ and $\hat{B}_{k+3}=B_{k+3},\ldots,\hat{B}_{m+2}=B_{m+2}$, i.e., that the factors for $\hat{Q}_{k+3}$,…, $\hat{Q}_{m+2}$ are unchanged. For $k+3$ we have
(19) $\hat{A}_{k+3}$ $=$ $\hat{a}_{k+1,1}Q_{1}+\cdots+\hat{a}_{k+1,k+1}Q_{k+1}+a_{k+1,k+2}\hat{Q}_{k+2}$
@@
where we have inserted (16a) and (18b) to obtain the second equality. The relation $\hat{B}_{k+3}=B_{k+3}$ can be shown analogously using (16b) and (18b). By induction, the corresponding relation for any factor. Consequently, we have
(20) $\hat{Q}_{i+2}=\hat{A}_{i+2}\hat{B}_{i+2}=A_{i+2}B_{i+2}=Q_{i+2},\quad i=k+1,\ldots,m$
@@
The conclusion (17) follows from the same construction as in (19). ∎
@@ 3.2 Normalized forms and unreduced schemes
The theorems in the previous section can be applied to impose a certain structure on the triplet $(A,B,c)$ without (or with very little) loss of generality. For any triplet $(A,B,c)$ we can invoke Theorem 2 repeatedly. By setting $\alpha$ to the negation of the element in the first column, we obtain matrices $A$ and $B$ with a first column containing only zeros. The first column corresponds to the addition of a scaled identity, which is independent (constant) with respect to the input $X$. Note that this can be done for any triplet, and no generality (in the sense of parameterized polynomials) is lost by assuming the first column is zero.
Definition 3.
@@
Similarly, if we assume that the Hessenberg matrices $A$ and $B$ are unreduced [11, p. 381], i.e., the elements $a_{1,2},\ldots,a_{m,m+1}$ and $b_{1,2},\ldots,b_{m,m+1}$ are all nonzero, we can impose further structure by repeatedly applying Theorem 1 with scaling determined by the corresponding last nonzero element of the row. This process imposes a normalization on each row.
Definition 5.
A constant-free evaluation scheme satisfying $a_{1,2}=\cdots=a_{m,m+1}=b_{1,2}=\cdots=b_{m,m+1}=1$ is called a normalized evaluation scheme, and we call the triplet $(A,B,c)$ normalized.
Corollary 6.
Any evaluation scheme corresponding to $(A,B,c)$ where $A$ and $B$ are unreduced Hessenberg matrices is equivalent to a normalized evaluation scheme.
@@ 3.3 Further transformations
For normalized constant-free evaluation schemes, we have $a_{11}=b_{11}=0$ and $a_{12}=b_{12}=1$, meaning that the first multiplication always corresponds to squaring the input matrix, i.e.,
@@
This assumption can be made without loss of generality and will be used henceforth. This fact was already observed by Paterson and Stockmeyer [19, p. 61].
We now derive further transformations under a technical assumption. For the first two rows we assume that we have the structure of an unreduced constant-free triplet. [^1] This assumption is made mostly to simplify the derivation, and similar results hold, e.g., when $a_{23}=0$. This in turn is equivalent to assuming $a_{23}=b_{23}=1$.
Consider the following transformation based on modifying the second row. We add $\alpha$ and $\beta$ to the second element in the second row in $A$ and $B$, respectively, i.e.,
@@
If we enforce $\beta=-\alpha$, one term in the modification disappears, and we can compensate for the other term in a way such that $\hat{Q}_{5}=Q_{5},\ldots,\hat{Q}_{m}=Q_{m}$, i.e., they are unmodified. Consequently, the output polynomial is also unmodified.
Theorem 7.
[…]
@@
To prove (27a), we use relation (22). More precisely, we substitute $\beta=-\alpha$ into (22b) and obtain
@@
To prove (27b), we show that the factors for $Q_{3},\ldots,Q_{m+2}$ are unchanged, i.e., $\hat{A}_{3}=A_{3},\ldots,\hat{A}_{m+2}=A_{m+2}$ and $\hat{B}_{3}=B_{3},\ldots,\hat{B}_{m+2}=B_{m+2}$. For the first factor equality, we have
(29a) $\hat{A}_{5}$ $=$ $a_{32}Q_{2}+\hat{a}_{33}Q_{3}+a_{34}\hat{Q}_{4}$
(29b) $=$ $a_{32}Q_{2}+a_{33}Q_{3}-a_{34}sQ_{3}+a_{34}Q_{4}+a_{34}sQ_{3}=A_{5},$
where we have used (25a) and (27a) with $i=3$, in the second equality. The relation $\hat{B}_{5}=B_{5}$ can be shown analogously using (25b) and (27a). By induction, we can prove the corresponding factor relation for $i=4,\ldots,m$. Consequently, we have $\hat{Q}_{i+2}=\hat{A}_{i+2}\hat{B}_{i+2}=A_{i+2}B_{i+2}=Q_{i+2}$ for $i=3,\ldots,m,$ which proves (27b). The theorem conclusion (26) follows by the same construction as (29). ∎
For a given evaluation scheme satisfying $a_{23}=b_{23}=1$, we can apply the previous theorem with $\alpha=b_{22}$, resulting in $\hat{b}_{22}=0$. This shows that we can assume $b_{22}=0$ for evaluation schemes that are unreduced in the first two rows, without loss of generality. For the second multiplication, under these assumptions, we get
@@
Next, we state and prove a theorem for the third row of the coefficient matrices. For this we assume that the first three rows are unreduced, i.e., $a_{23}=a_{34}=b_{23}=b_{34}=1$. The theorem is based on perturbing each of the free elements in the third row of $A$ and $B$, while simultaneously preserving the final output polynomial. In particular, we use perturbations with the following structure:
@@
It turns out that when the perturbations have this particular structure, $Q_{5}$ is modified by the addition of a linear combination of $Q_{3}$, $Q_{4}$ and $X^{3}$. This is advantageous because we can compensate for $Q_{3}$ and $Q_{4}$, since we have access to these matrices directly. Moreover, we can ensure that the $X^{3}$-coefficient modification is zero by placing an additional condition on $\alpha$ and $\beta$, encoded in the equality $z(\alpha,\beta)=0$, where $z$ is a function explicitly given in the theorem.
Theorem 8.
[…]
@@
where
(35a) $s_{1}$ $=$ $\alpha(b_{32}-a_{32})-\alpha^{2},$
@@
To prove (38a) we express $\hat{Q}_{5}$ in terms of the multiplication factors and substitute the definition of the modified coefficients. We obtain
@@
This expression can be simplified by using $Q_{5}=A_{5}B_{5}$:
(40) $\begin{split}\hat{Q}_{5}&=Q_{5}+(B_{5}-A_{5})(\alpha Q_{2}+\beta Q_{3})-(\alpha Q_{2}+\beta Q_{3})^{2}.\end{split}$
The next step is to show that the difference between $Q_{5}$ and $\hat{Q}_{5}$ is a linear combination of $Q_{3}$, $Q_{4}$, and $X^{3}$, where the $X^{3}$-coefficient is given by $z(\alpha,\beta)$, which is zero by assumption. We factorize the expression in (40) to obtain
@@
(42) $\begin{split}d_{32}:=b_{32}-a_{32},\\ d_{33}:=b_{33}-a_{33}.\end{split}$
@@
We substitute this expression into (41) and simplify the first factor of the difference $\hat{Q}_{5}-Q_{5}$ such that
[…]
Next, we expand the final expression
[…]
@@
where we have used that $z(\alpha,\beta)=0$ in the last equality. This proves statement (38a).
To prove (38b) we show that the multiplication factors are unchanged, i.e., $\hat{A}_{6}=A_{6},\ldots\hat{A}_{m+2}=A_{m+2}$ and $\hat{B}_{6}=B_{6},\ldots\hat{B}_{m+2}=B_{m+2}$. For the first factor we have
(48) $\begin{split}\hat{A}_{6}&=(a_{42}Q_{2}+\hat{a}_{43}Q_{3}+\hat{a}_{44}Q_{4}+a_{45}\hat{Q}_{5})\\ &=a_{42}Q_{2}+(a_{43}-a_{45}s_{1})Q_{3}+(a_{44}-a_{45}s_{2})Q_{4}+a_{45}Q_{5}+a_{45}(s_{1}Q_{3}+s_{2}Q_{4})\\ &=A_{6}-a_{45}s_{1}Q_{2}-a_{45}s_{2}Q_{4}+a_{45}(s_{1}Q_{3}+s_{2}Q_{4})\\ &=A_{6},\end{split}$
where we have used (34a), (34b) and (38a) in the second equality. The relation $\hat{B}_{6}=B_{6}$ follows analogously using (34c), (34d) and (38a).
@@
(49) $\hat{Q}_{i+2}=\hat{A}_{i+2}\hat{B}_{i+2}=A_{i+2}B_{i+2}=Q_{i+2},\quad i=4,\ldots,m,$
which proves (38b). The theorem conclusion (36) follows by the same construction as (48). ∎
Free variables in Theorem 8
Note that Theorem 8 includes a scalar-valued condition, $z(\alpha,\beta)=0$, involving two scalar variables. Let $d_{32}=b_{32}-a_{32}$ and $d_{33}=b_{33}-a_{33}$ be defined as in the proof of the theorem. If we let $\alpha$ be given, we get a quadratic equation in $\beta$:
@@
When $a_{22}\neq 0$, the solution $\beta$ to the equation is explicitly available from the solution of the quadratic equation. This has a disadvantage of introducing a square root operation. In a real setting, this can yield complex coefficients.
This is related to the result in paper [24] as follows. The results suggest several approaches to evaluate polynomials with a low number multiplications. For the case $m=3$, the formulas [24, Eq. (31)] involve a square root, and indeed that approach can be derived from the above transformation with $\alpha=-a_{32}$ and solving for $\beta$.
Suppose $\beta$ is given. Then, the solution for $\alpha$ can be expressed as
@@
with the condition $\beta\neq\frac{1}{2}d_{33}$. To reframe this condition in terms of entries in matrices $A$ and $B$, we use a change of variables that essentially generalizes [21] (where it is given for $m=3$):
@@
This leads to the relation
@@
Consequently, any evaluation scheme that is unreduced in the first three rows can be transformed into an equivalent scheme satisfying
@@
since we can choose any nonzero $r$. Thus, when considering evaluation schemes that are unreduced in the first three rows, we can assume without loss of generality that $b_{33}=a_{33}+1$. This assumption allows us to reduce the number of variables without introducing additional constraints, as explained in the following section.